OpenStudy (anonymous):

Find the general solution of the differential equation:

OpenStudy (anonymous):

\[\frac{ dr }{ dp }=4\sin p\]

OpenStudy (anonymous):

r=

OpenStudy (kainui):

So what you can do is play with the differential as though it were a fraction and multiply both sides by "dp" to get: dr=4sinp dp Now integrate and don't forget +C

OpenStudy (anonymous):

I must be doing something wrong I'm not getting the right answer and not sure where I'm going wrong

OpenStudy (kainui):

What's the right answer supposed to be?

OpenStudy (anonymous):

I don't know, but nothing I've tried is right

OpenStudy (kainui):

Show me what you get when you solve the integral I just wrote out above.

OpenStudy (anonymous):

I don't see an integral

OpenStudy (kainui):

dr=4sinp dp That is an indefinite integral without its fancy looking "S" shape.

OpenStudy (anonymous):

so the dr stands for the "S"?

OpenStudy (anonymous):

Ok i figured it out thanks

OpenStudy (kainui):

No, dr stands for infinitesimally small. What level of calculus are you in?

OpenStudy (anonymous):

Given \(dr=4\sin p\;dp,\) you can antidifferentiate both sides to get \(\int dr=\int 4\sin p\;dp.\)